### v1

@@ 3.2 Method 2: Dimension Reduction
where $\hat{\boldsymbol{n}}_{\partial S}$ is the conormal vector along $\partial S$. That is, $\hat{\boldsymbol{n}}_{\partial S}$ is tangential to $S$ and normal to $\partial S$. It is useful to note that we could use any boundary condition for (25). For the meshfree methods considered in this paper, we do not require any boundary conditions at all. According to Proposition 2.1 and Corollary 2.3, as long as (25) admits at least one solution in $\mathcal{H}$, a suitable symmetric meshfree method will produce solutions that converge to the norm-minimizing solution to (25) in $\mathcal{H}$. This emphasizes a key feature of these sorts of meshfree methods: problems that are ill-posed by virtue of admitting multiple solutions become well-posed when the norm-minimization condition is imposed.
Once the divergence theorem is applied to obtain (26), we are left with an integral on a lower-dimensional manifold. The process can then be repeated until only a line integral remains. The discretized problem for (25) is, for a point cloud $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}\subset S$,
@@
Let $u^{\left(\tilde{N}\right)}$ be the solution to this problem. If $f,S$, and the functions in $\mathcal{H}$ are sufficiently smooth, then according to (12), we have
@@
$\left|\int_{S}f-\int_{S}\Delta_{S}u^{\left(\tilde{N}\right)}\right|=\left|\int_{S}f-\int_{\partial S}\nabla_{S}u^{\left(\tilde{N}\right)}\cdot\hat{\boldsymbol{n}}_{\partial S}\right|$ $\leq A_{p}\left|S\right|h_{\text{max}}^{p}\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$,
where $\left|S\right|:=\int_{S}1$. This can be repeated for each dimension until $\int_{\partial S}\nabla_{S}u^{\left(\tilde{N}\right)}\cdot\hat{\boldsymbol{n}}_{\partial S}$ is simply a line integral.
@@ 3.3 Line Integrals
where $C$ is an open curve of interest and $\hat{\boldsymbol{t}}_{C}$ denotes its unit tangent vector. If we have a closed curve, such as the boundary of a surface, we first divide it into two open curves. Then, $\int_{C}f=u\left(\boldsymbol{b}\right)-u\left(\boldsymbol{a}\right)$, where $\boldsymbol{a}$ and $\boldsymbol{b}$ are the endpoints of $C$. To discretize (29), we use
@@
where $\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}$ is now a point cloud on $C$. Once again, with suitable smoothness assumptions, it is straightforward to show that
@@
for some $A_{p}>0$, where $u^{\left(\tilde{N}\right)}$ is the solution to (30) and $p$ depends on the smoothness of $f,S$, and the functions in $\mathcal{H}$.
@@ 3.4 Solving Surface Poisson Problems
A complication of solving (19) or (27) on surfaces is that it is not immediately obvious how to compute $\Delta_{S}u$ on a surface; $u$ is a function on a domain $\Omega\supset S$, not on the surface itself, and $\Delta u\biggr|_{S}\neq\Delta_{S}u$ in general. To address this, we note (see Lemma 1 of [34])
@@
We note that it is unnecessary to impose both $\hat{\boldsymbol{n}}_{S}\cdot\nabla u\left(\boldsymbol{x}_{j}\right)=0$ and $\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}=0$ here, although this has been done before in other surface PDE methods [20] and would work here as well. However, imposing both these conditions results in larger matrices, a longer computation time, and a smaller space of feasible interpolants. A smaller constraint set then leads to a larger value of $\left\|\tilde{u}\right\|_{\mathcal{H}}$ for the interpolant solution, which in turn can result in a less accurate solution. See [31] for more discussion on solving surface PDEs with this approach.
@@ 3.5.2 Planar Boundary
A simple approach to divide closed surfaces into two subdomains is to use a single plane defined by $P=\left\{\boldsymbol{x}\in\mathbb{R}^{3}:\boldsymbol{p}\cdot\boldsymbol{x}-c=0\right\}$ for some constant $c$. In this case, the two subdomains are given by
@@
and the boundaries $\partial S_{\pm}$ are closed planar curves. A complication that can arise for surfaces of non-zero genus is that each boundary may consist of a collection of disjoint curves. We note here that the meshfree approach is primarily advantageous for surfaces where high-order meshing may be challenging; integrals on curves may be computed more simply by other means. A straightforward technique for identifying and integrating over multiple planar, closed curves is as follows:
@@
    For each point in $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$, find the nearest point $\boldsymbol{y}_{j}^{\left(2\right)}$ such that $\left(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{\left(1\right)}\right)\cdot\left(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{\left(2\right)}\right)<0$ (so that $\boldsymbol{y}_{j}^{\left(1\right)}$ and $\boldsymbol{y}_{j}^{\left(2\right)}$ are on opposite sides of $\boldsymbol{y}_{j}$). That is, $\boldsymbol{y}_{j}^{\left(2\right)}:=\text{argmin}\left\{\left\|\boldsymbol{y}-\boldsymbol{y}_{j}\right\|:\boldsymbol{y}\in\left\{\boldsymbol{y}_{k}\right\}_{k=1}^{\tilde{N}_{\partial}},\left(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{\left(1\right)}\right)\cdot\left(\boldsymbol{y}_{j}-\boldsymbol{y}\right)<0\right\}$.
@@
    For each trio $\boldsymbol{y}_{j},\boldsymbol{y}_{j}^{\left(1\right)},\boldsymbol{y}_{j}^{\left(2\right)}$, construct the circle or line passing through each of the three points. Parametrize circles by the angle $\theta$ between the points on the circle, the circle’s centre, and $\boldsymbol{y}_{j}$. Parametrize lines by $s$ where $s\left(\boldsymbol{x}\right)=\left(\boldsymbol{x}-\boldsymbol{y}_{j}\right)\cdot\left(\boldsymbol{y}_{j}^{\left(1\right)}-\boldsymbol{y}_{j}\right)/\left\|\boldsymbol{y}_{j}^{\left(1\right)}-\boldsymbol{y}_{j}\right\|_{2}$ (signed distance from $\boldsymbol{y}_{j}$).
@@
This approach approximates the curve geometry by fitting a circle to each set of three points; approximating the curve geometry is necessary to have an order of convergence greater than two; this approach is third-order. Note that this method also naturally handles multiple curves by only using nearest neighbours. If $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$ is dense enough on each constituent curve of $\partial S_{\pm}$, each trio of neighbouring points will belong to the same closed curve, correctly approximating the geometry of $\partial S_{\pm}$.
@@ 4.1 Genus-Two Surface
(34) $\varphi\left(x,y,z\right):=\frac{1}{4\left(\left(x-1\right)^{2}+y^{2}\right)}+\frac{1}{4\left(\left(x+1\right)^{2}+y^{2}\right)}+\frac{1}{10}x^{2}+\frac{1}{4}y^{2}+z^{2}-1.$
@@ 4.1.1 Average Values
Recall that Method 1 computes ratios of integrals: $\int_{S}f/\int_{S}g$. This means that it is straightforward to compute average values by setting $g=1$. To start, we use approximately uniformly generated point clouds using version B of algorithm A.1 of [30] with 25 test points per point in $S_{\tilde{N}}$. From a highly refined computation, we find that the average value of $x^{2}$ on the surface given by the zero set of $\varphi$ from (34) is 2.45884. We compare Method 1, using (33) to solve the Laplace-Beltrami problem, to integrating on a triangle mesh generated with the ball-and-pivot method [1] as implemented by the MATLAB Lidar Toolbox’s [27] $\mathtt{pc2surfacemesh}$, using linear interpolation between vertices (as in P1 finite elements). Integrating over the triangulation using a higher-order interpolation method would require a high-order approximation of the surface geometry and interpolation of additional points inside or outside the triangle (see [22] for example, for constructing a more accurate interpolant on a triangulation using RBFs).
@@
(35) $d_{x,n_{x}}=\left(\exp\left(q\sqrt{\frac{2\pi}{T}}\right)+\exp\left(q\sqrt{\left|\omega_{x,n_{x}}\right|}\right)\right)^{2}.$
$d_{y,n_{y}}$ and $d_{z,n_{z}}$ are defined similarly, substituting $\omega_{y,n_{y}}$ and $\omega_{z,n_{z}}$, respectively, in place of $\omega_{x,n_{x}}$. We use $T=12$, $\Omega=\left[-5,5\right]^{3}$ and $\left(2\cdot 80+1\right)^{3}$ Fourier basis functions (treated separably to form the $\boldsymbol{\Phi}$ matrix so that this is computationally feasible) to generate Figure 1, which shows the convergence of the meshfree method compared against triangulation and simply taking the average value on the point cloud (Point Cloud Average), as in a Monte Carlo method. The reference value we use to estimate the relative error is computed using a triangulation with 4.8 million vertices. We test a couple of different values of $q$; our observation is that for domains roughly of this size, a choice of $q$ between 2 and 4 works reasonably well. Functions with more oscillations are often better approximated using a lower value of $q$ to reduce the suppression of higher frequencies. However, as demonstrated in Figure 1, an optimal selection of $q$ is not necessary to achieve rapid convergence, and multiple values of $q$ could work well. It should be noted that the points are not quite uniformly generated, which explains why the “Point Cloud Average” method fails to converge. The point cloud with $\tilde{N}=800$ is shown in Figure 2 (Left).
@@
There are a few observations to make regarding Figure 1. For the triangulation approach, the error is high before falling into a pattern of consistent $\mathcal{O}\left(h_{\text{max}}^{2}\right)=\mathcal{O}\left(\tilde{N}^{-1}\right)$ convergence; this is due to meshing errors at low point resolutions. The meshfree methods reach a relative error on the order of $10^{-5}$ using only around two or three thousand points. This is roughly the error floor using the $\boldsymbol{\Phi}$ matrix, so other strategies (namely, complete orthogonal decomposition or singular value decomposition) would have to be employed to solve the optimization problem with higher accuracy. For a fast and reasonably accurate approach using a sparse point cloud, however, the separable basis method excels. It would take a mesh with nearly 100000 vertices to reach the level of accuracy achieved by the meshfree method with under 3000 points.
Also of note is that the meshfree approach does not require the point cloud to be “well-spaced”. That is, the point cloud may be significantly denser in some regions of the surface or have completely random point spacing, and the meshfree approach will still produce a reasonable result. Monte Carlo methods will clearly fail when points are drawn from an unknown, non-uniform distribution, and meshing algorithms tend to produce errors when the point cloud spacing is extremely uneven. We use randomly generated point clouds using algorithm A.1 of [30] (1 test point per point) for the tests in Table 1 and use $q=3$ for the meshfree method, with the other parameters left unchanged from Figure 1. Unlike in the previous test, the point cloud is not processed at all here after points are placed on the surface; the points are not as evenly spaced as in the previous test. This is shown in Figure 2 (Centre). Meshing these point clouds using MATLAB’s ball-and-pivot routine produces many missing elements, so we do not include it in this test.
@@
Table 1: Left: Relative error for computing the average value of $x^{2}$ on a genus-two surface and using randomly generated points (Figure 2, Centre) and an uneven point distribution (Figure 2, Right). We use $q=3$ for both tests.

### v2

@@ 3.2 Method 2: Dimension Reduction
where $\hat{\boldsymbol{n}}_{\partial S}$ is again the conormal vector along $\partial S$. It is useful to note that we could use any boundary condition for (25). For the meshfree methods considered in this paper, we do not require any boundary conditions at all. According to Proposition 2.1 and Corollary 2.3, as long as (25) admits at least one solution in $\mathcal{H}$, a suitable symmetric meshfree method will produce solutions that converge to the norm-minimizing solution to (25) in $\mathcal{H}$. This emphasizes a key feature of these sorts of meshfree methods: problems that are ill-posed by virtue of admitting multiple solutions become well-posed when the norm-minimization condition is imposed.
Once the divergence theorem is applied to obtain (26), we are left with an integral on a lower-dimensional manifold. The process can then be repeated until only a line integral remains. The discretized problem for (25) is, for a point cloud $\{\boldsymbol{x}_{j}\}_{j=1}^{\tilde{N}}\subset S$,
@@
Let $u^{(\tilde{N})}$ be the solution to this problem. If $f,S$, and the functions in $\mathcal{H}$ are sufficiently smooth, then according to (12), we have
@@
$\left|\int_{S}\Delta_{S}u^{\left(\tilde{N}\right)}-\int_{S}f\right|=\left|\int_{\partial S}\nabla_{S}u^{\left(\tilde{N}\right)}\cdot\hat{\boldsymbol{n}}_{\partial S}-\int_{S}f\right|$ $\leq A_{p}\left|S\right|h_{\text{max}}^{p}\left\|u^{\left(\tilde{N}\right)}\right\|_{\mathcal{H}}$,
where $\left|S\right|:=\int_{S}1$. This can be repeated for each dimension until $\int_{\partial S}\nabla_{S}u^{(\tilde{N})}\cdot\hat{\boldsymbol{n}}_{\partial S}$ is simply a line integral.
@@ 3.3 Line Integrals
where $C$ is an open curve of interest and $\hat{\boldsymbol{t}}_{C}$ denotes its unit tangent vector. If we have a closed curve, such as the boundary of a surface, we first divide it into two open curves. Then, $\int_{C}f=u\,(\boldsymbol{b})-u\,(\boldsymbol{a})$, where $\boldsymbol{a}$ and $\boldsymbol{b}$ are the endpoints of $C$. To discretize (29), we use
@@
where $\{\boldsymbol{x}_{j}\}_{j=1}^{\tilde{N}}$ is now a point cloud on $C$. Once again, with suitable smoothness assumptions, it is straightforward to show that
@@
for some $A_{p}>0$, where $u^{(\tilde{N})}$ is the solution to (30) and $p$ depends on the smoothness of $f$, $C$, and the functions in $\mathcal{H}$.
@@ 3.4 Solving Surface Poisson Problems
A complication of solving (19) or (27) on surfaces is that it is not immediately obvious how to compute $\Delta_{S}u$ on a surface; $u$ is a function on a domain $\Omega\supset S$, not on the surface itself, and $\Delta u|_{S}\neq\Delta_{S}u$ in general. To address this, we note (see Lemma 1 of [39])
@@
We note that it is unnecessary to impose both $\hat{\boldsymbol{n}}_{S}\cdot\nabla u\left(\boldsymbol{x}_{j}\right)=0$ and $\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}u\right)\hat{\boldsymbol{n}}_{S}=0$ here, although this has been done before in other surface PDE methods [24] and would work here as well. However, imposing both these conditions results in larger matrices, a longer computation time, and a smaller space of feasible interpolants. A smaller constraint set then leads to a larger value of $\|\tilde{u}\|_{\mathcal{H}}$ for the interpolant solution, which in turn can result in a less accurate solution. See [36] for more discussion on solving surface PDEs with this approach.
@@ 3.5.2 Planar Boundary
A simple approach to divide closed surfaces into two subdomains is to use a single plane defined by $P=\{\boldsymbol{x}\in\mathbb{R}^{3}:\boldsymbol{p}\cdot\boldsymbol{x}-c=0\}$ for some constant $c$. In this case, as long as $P$ is chosen such that it is nowhere tangent to the surface, the two subdomains are given by
@@
and the boundaries $\partial S_{\pm}$ are closed planar curves. A complication that can arise for closed surfaces that bound a non-convex volume is that each boundary may consist of a collection of disjoint curves. We note here that the meshfree approach is primarily advantageous for surfaces where high-order meshing may be challenging; integrals on curves may be computed more simply by other means. A straightforward technique for identifying and integrating over multiple planar, closed curves is as follows:
@@
    For each point in $\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}$, find the nearest point $\boldsymbol{y}_{j}^{\left(2\right)}$ such that $(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{(1)})\cdot(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{(2)})<0$ (so that $\boldsymbol{y}_{j}^{\left(1\right)}$ and $\boldsymbol{y}_{j}^{\left(2\right)}$ are on opposite sides of $\boldsymbol{y}_{j}$). That is, $\boldsymbol{y}_{j}^{\left(2\right)}:=\text{argmin}\,\{\|\boldsymbol{y}-\boldsymbol{y}_{j}\|_{2}:\boldsymbol{y}\in\{\boldsymbol{y}_{k}\}_{k=1}^{\tilde{N}_{\partial}},(\boldsymbol{y}_{j}-\boldsymbol{y}_{j}^{(1)})\cdot\left(\boldsymbol{y}_{j}-\boldsymbol{y}\right)<0\}$.
@@
    For each trio $\boldsymbol{y}_{j},\boldsymbol{y}_{j}^{\left(1\right)},\boldsymbol{y}_{j}^{\left(2\right)}$, construct the circle or line passing through each of the three points. Parametrize circles by the angle $\theta$ between the points on the circle, the circle’s centre, and $\boldsymbol{y}_{j}$. Parametrize lines by $s$ where $s\left(\boldsymbol{x}\right)=(\boldsymbol{x}-\boldsymbol{y}_{j})\cdot(\boldsymbol{y}_{j}^{(1)}-\boldsymbol{y}_{j})/\|\boldsymbol{y}_{j}^{\left(1\right)}-\boldsymbol{y}_{j}\|_{2}$ (signed distance from $\boldsymbol{y}_{j}$).
@@
This approach approximates the curve geometry by fitting a circle to each set of three points; approximating the curve geometry is necessary to have an order of convergence greater than two; this approach is third-order. Note that this method also naturally handles multiple curves by only using nearest neighbours. If $\{\boldsymbol{y}_{j}\}_{j=1}^{\tilde{N}_{\partial}}$ is dense enough on each constituent curve of $\partial S_{\pm}$, each trio of neighbouring points will belong to the same closed curve, correctly approximating the geometry of $\partial S_{\pm}$.
@@ 4.1 Genus-Two Surface
(34) $\varphi\left(x,y,z\right):=\frac{1}{4}\left(\left(x-1\right)^{2}+y^{2}\right)^{-1}+\frac{1}{4}\left(\left(x+1\right)^{2}+y^{2}\right)^{-1}+\frac{1}{10}x^{2}+\frac{1}{4}y^{2}+z^{2}-1.$
@@ 4.1.1 Average Values
Recall that Method 1 computes ratios of integrals: $\int_{S}f/\int_{S}g$. This means that it is straightforward to compute average values by setting $g=1$. To start, we use approximately uniformly generated point clouds using version B of algorithm A.1 of [35] with 25 test points per point in the surface point cloud $S_{\tilde{N}}$. From a highly refined computation, we find that the average value of $x^{2}$ on the surface given by the zero set of $\varphi$ from (34) is 2.45884. We compare Method 1, using (33) to solve the Laplace-Beltrami problem, to integrating on a triangle mesh generated with the ball-and-pivot method [2] as implemented by the MATLAB Lidar Toolbox’s [31] $\mathtt{pc2surfacemesh}$, using linear interpolation between vertices (as in P1 finite elements). Integrating over the triangulation using a higher-order interpolation method would require a high-order approximation of the surface geometry and interpolation of additional points inside or outside the triangle (see [27] for example, for constructing a more accurate interpolant on a triangulation using RBFs).
@@
(35) $d_{x,n_{x}}=\left(\exp\left(q\sqrt{2\pi/T}\right)+\exp\left(q\sqrt{\left|\omega_{x,n_{x}}\right|}\right)\right)^{2}.$
$d_{y,n_{y}}$ and $d_{z,n_{z}}$ are defined similarly, substituting $\omega_{y,n_{y}}$ and $\omega_{z,n_{z}}$, respectively, in place of $\omega_{x,n_{x}}$. We use $T=12$, $\Omega=[-5,5]^{3}$ and $(2\cdot 80+1)^{3}$ Fourier basis functions (treated separably to form the $\boldsymbol{\Phi}$ matrix so that this is computationally feasible) to generate Figure 1, which shows the convergence of the meshfree method compared against triangulation and simply taking the average value on the point cloud (Point Cloud Average), as in a Monte Carlo method. The reference value we use to estimate the relative error is computed using a triangulation with 4.8 million vertices.
We test a couple of different values of $q$; our observation is that for domains roughly of this size, a choice of $q$ between 2 and 4 works reasonably well. Functions with more oscillations are often better approximated using lower values of $q$ and $T$ to reduce the suppression of higher frequencies. However, as demonstrated in Figure 1, an optimal selection of $q$ is not necessary to achieve rapid convergence, and multiple values of $q$ (or $T$) could work well. The number of Fourier basis functions should be substantially larger than the total number of interpolation conditions, and larger still to capture rapid oscillations; a more detailed discussion of the selection of the number of frequencies for solving PDEs is in Remark 2 and Proposition 9 of [36]. It should be noted that the points are not quite uniformly generated, which explains why the “Point Cloud Average” method fails to converge. The point cloud with $\tilde{N}=800$ is shown in Figure 2 (Left).
@@
There are a few observations to make regarding Figure 1. For the triangulation approach, the error is high before falling into a pattern of consistent $\mathcal{O}\,(h_{\text{max}}^{2})=\mathcal{O}\,(\tilde{N}^{-1})$ convergence; this is due to meshing errors at low point resolutions. The meshfree methods reach a relative error on the order of $10^{-5}$ using only around two or three thousand points. This is roughly the error floor using the $\boldsymbol{\Phi}$ matrix, so other strategies (namely, complete orthogonal decomposition or singular value decomposition) would have to be employed to solve the optimization problem with higher accuracy. For a fast and reasonably accurate approach using a sparse point cloud, however, the separable basis method excels. It would take a mesh with nearly 100000 vertices to reach the level of accuracy achieved by the meshfree method with under 3000 points.
We recall that the first $\tilde{N}$ components of $\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}/(\boldsymbol{g}^{*}\boldsymbol{\Phi}^{-1}\boldsymbol{g})$ can be interpreted as integration weights for computing average values; call the vector consisting of these components $\tilde{\boldsymbol{w}}^{*}$. While not a factor in the error estimates of 3.1.1 as it may be in integration schemes based on exactly integrating an interpolant (see, for example, the discussion around Eq. (2.4) of [34]), the quantity $\kappa:=\sum_{j=1}^{\tilde{N}}|\tilde{w}_{j}|/\sum_{j=1}^{\tilde{N}}\tilde{w}_{j}$, which is equal to $\left\|\tilde{\boldsymbol{w}}\right\|_{1}$ when $g=1$, is still relevant for stability. Specifically, for two functions $f_{1},f_{2}$, our average value estimate will differ by at most $\|\tilde{\boldsymbol{w}}\|_{1}\|f_{1}-f_{2}\|_{L^{\infty}(S)}$. This means the method’s ability to handle noise, rounding errors, and small deviations from smooth functions is determined by $\left\|\tilde{\boldsymbol{w}}\right\|_{1}$. There is naturally a trade-off between accuracy and stability; this is demonstrated in Table 1, which shows $\|\tilde{\boldsymbol{w}}\|_{1}$ and the error in the average value of $x^{2}$ on $S$ for select values of $\tilde{N},q,$ and $T$. We see that all values of $\|\tilde{\boldsymbol{w}}\|_{1}$ remain fairly close to one, but higher values of $T$ or $q$ generally produce more accurate results at the cost of a higher value of $\|\tilde{\boldsymbol{w}}\|_{1}$.
$\tilde{N}=1600$ $\tilde{N}=2400$
$(q,T)$ Relative Error $\left\|\tilde{\boldsymbol{w}}\right\|_{1}$ Relative Error $\left\|\tilde{\boldsymbol{w}}\right\|_{1}$
(7/3, 2) 7.2741E-04 1.03 1.3862E-04 1.05
(7/3, 12) 3.8281E-04 1.10 7.9380E-05 1.14
(10/3, 2) 1.8272E-05 1.08 9.7313E-06 1.12
(10/3, 12) 3.6092E-06 1.19 1.1916E-05 1.26
Table 1: Relative error for computing $x^{2}$ on a genus-two surface and $\left\|\tilde{\boldsymbol{w}}\right\|_{1}$ for various combinations of $\tilde{N},q$, and $T$.
An important note is that the meshfree approach does not require the point cloud to be “well-spaced”. That is, the point cloud may be significantly denser in some regions of the surface or have completely random point spacing, and the meshfree approach will still produce a reasonable result. Monte Carlo methods will clearly fail when points are drawn from an unknown, non-uniform distribution, and meshing algorithms tend to produce errors when the point cloud spacing is extremely uneven. We use randomly generated point clouds using algorithm A.1 of [35] (1 test point per point) for the tests in Table 2 and use $q=3$ for the meshfree method, with the other parameters left unchanged from Figure 1. Unlike in the previous test, the point cloud is not processed at all here after points are placed on the surface; the points are not as evenly spaced as in the previous test. This is shown in Figure 2 (Centre). Meshing these point clouds using MATLAB’s ball-and-pivot routine produces many missing elements, so we do not include it in this test.
@@
Table 2: Left: Relative error for computing the average value of $x^{2}$ on a genus-two surface, using randomly generated points (Figure 2, Centre) and an irregular point distribution (Figure 2, Right). We use $q=3$ for both tests.